# Diffusion-empowered AutoPrompt MedSAM

### v1

@@ Abstract
MedSAM, a medical foundation model derived from the SAM architecture, has demonstrated notable success across diverse medical domains. However, its clinical application faces two major challenges: the dependency on labor-intensive manual prompt generation, which imposes a significant burden on clinicians, and the absence of semantic labeling in the generated segmentation masks for organs or lesions, limiting its practicality for non-expert users. To address these limitations, we propose AutoMedSAM, an end-to-end framework derived from SAM, designed to enhance usability and segmentation performance. AutoMedSAM retains MedSAM’s image encoder and mask decoder structure while introducing a novel diffusion-based class prompt encoder. The diffusion-based encoder employs a dual-decoder structure to collaboratively generate prompt embeddings guided by sparse and dense prompt definitions. These embeddings enhance the model’s ability to understand and process clinical imagery autonomously. With this encoder, AutoMedSAM leverages class prompts to embed semantic information into the model’s predictions, transforming MedSAM’s semi-automated pipeline into a fully automated workflow. Furthermore, AutoMedSAM employs an uncertainty-aware joint optimization strategy during training to effectively inherit MedSAM’s pre-trained knowledge while improving generalization by integrating multiple loss functions. Experimental results across diverse datasets demonstrate that AutoMedSAM achieves superior performance while broadening its applicability to both clinical settings and non-expert users. Code is available at https://github.com/HP-ML/AutoPromptMedSAM.git.
@@ 1 Introduction
Deep learning models have traditionally been applied in medicine by designing and training specialized models for specific tasks, achieving significant success [1, 2, 3, 4]. However, these approaches often require training model from scratch using corresponding data [5, 6, 7], leading to low training efficiency and limited transferability to other tasks [8, 9]. Recently, prompt-based foundational models in computer vision, such as the Segment Anything Model (SAM) [10], have demonstrated impressive performance and generalization capabilities in various semantic segmentation tasks based on user-provided prompts [11, 12].
@@
Although MedSAM boasts impressive zero-shot and few-shot segmentation capabilities, it is essentially a fine-tuned version of SAM without changes to its prompt-based framework. This means that MedSAM retains both the advantages and shortcomings of SAM [14, 10]. Specifically: (1) Dependency on Manual Prompts: It requires users to provide precise prompts to segment the target region. However, creating prompts for medical images requires expertise, and in clinical settings, providing explicit prompts like points, bounding boxes, or scribbles is impractical. (2) Prompt Encoder Limitations: The Prompt Encoder limits MedSAM’s segmentation capabilities. The manual nature of these prompts means the error may not be within a controllable range [15] Using the bounding box prompt as an example, MedSAM’s performance relies on the discrepancy between the prompt box and the true boundary. However, other targets within the prompt box cannot be avoided, and different organ or lesion categories often exhibit high similarity and low inter-class variation [16]. (3) Lack of Semantic Information: MedSAM cannot obtain the semantic information of the masks it predicts. It can only predict binary masks for each prompt, without associating them with semantic labels [10, 13, 17].
To address these challenges, we propose AutoMedSAM. A comparison between AutoMedSAM and existing methods is shown in Fig. 1. We replaced the original prompt encoder with a diffusion-based class prompt encoder. This new encoder uses lesion or organ index as a prompt, incorporating semantic information about the target into the AutoMedSAM learning process. The class prompts and image embeddings are input information to generate prompt embeddings for the mask decoder directly. By generating prompt embeddings directly from class prompts, we eliminate the robustness issues caused by manually provided explicit prompts and transform the semi-automatic MedSAM into a fully automated end-to-end process. Meanwhile, we designed an uncertainty-aware joint optimization training strategy. This strategy can optimize the model training process by combining the advantages of multiple loss functions while transferring the pre-training knowledge from MedSAM to AutoMedSAM. This enables AutoMedSAM to adapt to data of various modalities and effectively extract features of different organs or lesions, enhancing the segmentation performance and robustness of the model.
@@
    We introduced AutoMedSAM, which overcomes the limitations of the prompt-based SAM with a novel architecture, expanding the model’s applicability to a broader range of users and scenarios. This method is adaptable to more medical contexts and, through an end-to-end workflow, enables non-medical experts to obtain accurate segmentation results without relying on manual prompts.
@@
    We propose a diffusion-based class prompt encoder that eliminates the need for explicit prompts while adding semantic labels to the predicted masks. During the diffusion process, the class prompt is projected, allowing the model to capture class-specific features better. The unique two-decoder structure, combined with the definitions of sparse and dense prompt embeddings, enhances the controllability of the generation process.
@@
    We designed an uncertainty-aware joint optimization method that efficiently transfers the pre-trained knowledge from MedSAM to AutoMedSAM, while dynamically leveraging multiple loss functions to enhance the model’s generalization ability across multi-modal data.
@@ 2.1 SAM-based Medical Image Segmentation
SAM represents a significant breakthrough in transforming image segmentation from specialized applications to a general-purpose tool [18]. Its extensive pretraining on large-scale data has endowed it with remarkable zero-shot generalization capabilities [19]. Research on SAM-based medical image segmentation can be broadly divided into two categories. The first category focuses on fine-tuning while retaining the manual prompt structure to enhance performance. MedSAM was trained with a large volume of medical data to create a medical version of SAM [13]. MedficientSAM employs the EfficientViT model to replace the image encoder in SAM and improves the performance of the model after fine-tuning it on multimodal data [20]. However, manual prompts are difficult to provide accurately in clinical settings [16], leading to a second category of research focusing on replacing manual prompts to make SAM a fully automated process, as seen in models like MaskSAM and UV-SAM [17, 21]. These approaches introduce additional models to generate rough masks, but the added structures are often task-specific, making the SAM model more cumbersome. Adaptive SAM and SP-SAM utilize CLIP to encode textual prompts into prompt embeddings [22, 23]. Although these methods achieve a fully automated segmentation process, they still rely on prompts, leaving the issues outlined in Section  1 unresolved. Our proposed AutoMedSAM not only eliminates the need for manual prompts but also fully taps into SAM’s inherent potential. Moreover, the class prompt encoder can be extended to all SAM-based foundational models.
2.2 Diffusion Models for Medical
Diffusion models show strong potential in medical imaging, achieving notable success in tasks such as image generation, segmentation, and classification [24, 25, 26]. The initial applications of diffusion models in the medical field primarily focused on generating medical data, which has proven useful for medical data augmentation [27]. In addition, several scholars have investigated the potential of medical images generated by diffusion models as a substitute for real data in training deep networks. M. Usman Akbar et al. and D. Stojanovski et al. demonstrated that these synthetic data are effective for downstream tasks [28, 29]. Recent studies have used diffusion models for cross-modality synthesis [30, 31]. For example, DCE-diff addresses data heterogeneity by leveraging multimodal non-contrast images to extract anatomical details from structural MRI sequences and perfusion information from ADC images, enabling the synthesis of early- and late-phase DCE-MRI [32]. D. Stojanovski et al. demonstrated that visual realism during model training does not necessarily correlate with model performance. Consequently, these models can generate organ or lesion features optimized for deep learning, thereby enhancing the accuracy of downstream tasks. Furthermore, utilizing more efficient models can significantly reduce computational costs [29].
@@ 3.1 Overview of AutoMedSAM
The optimization of AutoMedSAM is essentially based on the structure of SAM, with the innovative introduction of a diffusion-based class prompt encoder to address the challenges of manual prompts. Specifically, as illustrated in Fig. 2, AutoMedSAM consists of three core modules: an image encoder $E_{I}$, a diffusion-based class prompt encoder $E_{P}$, and a mask decoder $D_{M}$. The input image is denoted as $I\in\mathbb{R}^{h\times w\times 3}$, with spatial resolution $h\times w$. By providing the prompt class $c$, AutoMedSAM can predict the mask corresponding to the class $c$. The image is first processed by the image encoder to generate image embedding $F_{I}$. Subsequently, the class prompt encoder $E_{P}$, based on a diffusion model, processes the image embedding to generate sparse prompt embedding $P_{s}^{(c)}$ and dense prompt embedding $P_{d}^{(c)}$ from the target prompts. Finally, the mask decoder combines the image embedding, positional encoding $P_{p}$, sparse prompt embedding, and dense prompt embedding to predict the segmentation mask $M^{(c)}$ for class $c$. The entire process can be represented as:
@@ 3.3 Uncertainty-aware Joint Optimization
AutoMedSAM is designed to accommodate various modalities of medical images, which often have significant differences. Moreover, medical images generally have characteristics such as low contrast and high similarity among targets. To enhance segmentation accuracy, the model’s optimization process needs to integrate multiple loss functions to improve overall performance. However, as the number of loss functions increases, adjusting their weights becomes increasingly challenging, and these weights are crucial for model optimization as they directly influence the optimization direction [35]. To address this issue, we introduced a mechanism that incorporates uncertainty-aware into the loss weights setting, combining pixel-based, region-based, and distribution-based loss functions to evaluate the generated prompt embeddings and segmentation masks. By fully utilizing the complementary advantages of these loss functions, we assess and optimize model performance comprehensively. Specifically, we employed the following loss functions:
@@
(3) Cross-Entropy (CE) Loss: A distribution-based loss that focuses on the difference between the predicted probability distribution and the actual class distribution.
@@
$L=\sum_{m=1}^{M}\left(\frac{1}{2\lambda_{m}^{2}}L_{m}+\log(1+\lambda_{m}^{2})\right),$ (17)
where $M$ represents the number of loss members involved in the optimization during training, $L_{m}$ represents all the losses we mentioned above, and $\lambda_{m}$ are learnable parameters that adjust the contribution of each loss component based on uncertainty.
@@ 4.1.1 Datasets
To evaluate the generalizability of AutoMedSAM, we conducted tests on several commonly used benchmark datasets, including AbdomenCT1K[37], BraTS[38], Kvasir-SEG[39], and Chest Xray Masks and Labels[40]. All data were obtained from the CVPR 2024 Medical Image Segmentation on Laptop Challenge [^1] https://www.codabench.org/competitions/1847. These datasets encompass four distinct imaging modalities: CT, MR, endoscopy, and X-ray, covering diverse segmentation targets ranging from organs to lesions. This diversity effectively demonstrates the broad applicability of our method. To equip the model for handling multimodal data, 3D images are converted into a three-channel format via slicing. Table 1 provides an overview of the datasets used.
@@
[…]
Method Category Method DSC(%$\uparrow$ NSD(%$\uparrow$ Organ
DSC(%$\uparrow$ NSD(%$\uparrow$
SAM-Core Model SAM[10] 89.79 83.94 92.117 93.766 72.694 95.254 95.121 75.376 90.253 69.687 92.385 91.998
SAM2[41] 90.191 85.137 93.815 95.391 70.793 95.623 95.334 81.804 94.149 62.519 93.86 93.354
[…]
Med2d[42] 83.84 79.347 93.562 93.371 58.168 87.081 87.02 80.02 89.751 67.776 79.624 79.566
SAM-Based model SAMed[44] 81.329 78.504 97.132 96.838 76.512 67.904 68.259 92.435 95.39 77.635 63.342 63.716
[…]
@@ 4.1.5 Baseline Methods
We compared our method with state-of-the-art medical imaging SAM models, grouped into SAM-Core models (SAM [10], SAM2 [41], MedSAM [13], Med2d [42], and U-MedSAM [43]) and SAM-Based models (SAMed [44], H-SAM [45], AutoSAM [46], and SurgicalSAM [16]). The SAM-Core Model retains the foundational framework of SAM, requiring manual prompting for operation. The SAM-Based Model builds upon SAM, introducing enhancements that eliminate the need for manual prompting. Notably, SurgicalSAM, like our approach, uses class prompts and prototype contrastive learning to distinguish surgical instruments. To ensure the reliability of the experimental results, all methods were conducted under identical experimental conditions.
@@ 4.2 Comparing with the Existing Methods
Prompt-based models should demonstrate the ability to identify multiple medical targets. Accordingly, we first evaluated the performance of various models on the multi-organ segmentation task using the AbdomenCT1K dataset. The experimental results are shown in Table 2. It is evident that the proposed AutoMedSAM significantly outperforms other models, achieving average DSC and NSD scores of 94.58% and 95.148%, respectively, which are improvements of 1.075% and 2.511% over the second-best MedSAM. Moreover, except for the spleen, AutoMedSAM achieved the best results across all organs, with DSC and NSD scores exceeding those of the lowest-performing SurgicalSAM by 19.075% and 25.029%, respectively. These findings demonstrate the clear superiority of AutoMedSAM in multi-organ segmentation for medical applications. Further analysis shows that SAM-Core models outperform SAM-Based models on average. Among SAM-Core models, even the lowest-performing Med2D outperforms all SAM-Based models except for our proposed model. This difference is mainly attributable to SAM-Core models employing manual prompt mechanisms, which provide precise positional information, enabling the model to focus more effectively and deliver stable performance.
@@
As shown in Table 2, the segmentation performance of the tested models varies minimally for the liver, spleen, and pancreas but significantly deteriorates for kidney segmentation. To investigate this, we visualized the segmentation results (Fig. 4). Subfigures (a) and (b) in Fig. 4 illustrate that while SAM-Based models generally perform well in boundary recognition, they confuse the left and right kidneys. This is why their performance drops sharply. However, during manual prompting, bounding boxes may unavoidably include other organs. Since SAM-Core models lack semantic information during prediction, they cannot accurately identify specific organs, leading to segmentation errors. As shown in subfigures (c) and (d) of Fig. 4, both SAM and MedSAM mistakenly identified other tissues to varying degrees. Additionally, even with minimal redundancy within bounding boxes, SAM models tend to segment features that appear prominent, which degrades mask quality (Fig. 4, subfigures (e) and (f)). In contrast, our method incorporates class-based prompts to introduce semantic information into predictions, effectively mitigating this issue.
@@
To evaluate segmentation accuracy independently, we conducted additional experiments on the BraTS, Kvasir-SEG, and Chest Xray Masks and Labels datasets. In BraTS and Kvasir-SEG, the targets (tumor and polyp) are singular but structurally complex, with diverse shapes and ambiguous boundaries. In Chest Xray Masks and Labels, overlapping structures like ribs around the lungs and the large target area introduce significant challenges. These datasets provide diverse difficulties suitable for evaluating segmentation performance. The results and visualizations are shown in Table 3 and Fig. 5, respectively.
Table 3: Comparative Results on the BraTS, Kvasir-SEG and Chest Xray Masks and Labels Datasets
@@
BraTS(Tumor) Kvasir-SEG(Polyp) Chest Xray Masks and Labels(Lung)
@@
U-MedSAM[43] 89.616 88.793 95.007 97.091 96.813 97.31
SAM-Based model SAMed[44] 89.902 89.097 86.393 88.255 94.066 94.6
H-SAM[45] 90.615 90.288 88.01 90.084 92.264 92.978
AutoSAM[46] 90.24 89.186 87.045 89.288 89.616 90.516
@@
[…]
As shown in Table 3, AutoMedSAM achieved superior performance across all tasks, with DSC and NSD scores of 96.828% and 98.729% for polyp segmentation, demonstrating its adaptability to complex medical environments. Conversely, when SAM-Core models lose their advantage of manual prompts, their performance declines significantly. For example, in BraTS, SAM achieved only 69.667% DSC and 42.112% NSD for tumor segmentation, indicating reasonable overlap with ground truth but poor boundary recognition. Similar issues were observed in lung segmentation, as shown in the last two subfigures of Fig. 5, where SAM models produced masks containing excessive non-target tissues. Finally, although SurgicalSAM also uses class-based prompts, its prototype-based contrastive learning framework reduces its adaptability in low-contrast, indistinct-boundary medical segmentation scenarios, limiting its ability to produce high-quality masks. In conclusion, the experimental results confirm that the proposed AutoMedSAM effectively adapts to various medical modalities, delivering accurate segmentation masks and achieving the best metrics across all datasets.
Figure 5: The qualitative analysis results of AutoMedSAM and other comparison models on BraTS, Kvasir-SEG, and Chest Xray Masks and Labels.

### v2

@@ Abstract
MedSAM, a medical foundation model derived from SAM, has achieved remarkable success across various medical domains. Nonetheless, its clinical adoption encounters two primary challenges: the reliance on labor-intensive manual prompt generation, which places a considerable burden on clinicians, and the lack of semantic labeling in the generated segmentation masks for organs or lesions, which hinders its usability for non-expert users. To overcome these limitations, we propose an end-to-end interactive segmentation model, AutoMedSAM. AutoMedSAM employs a diffusion-based class prompt encoder to generate prompt embeddings guided by the prompt class indices and eliminates the reliance on manual prompts. During the diffusion process, the encoder progressively captures the semantic structure and fine-grained features of the target object, injecting semantic information into the prediction pipeline. Furthermore, we introduce an uncertainty-aware joint optimization strategy that integrates pixel-based, region-based, and distribution-based losses. This approach harnesses MedSAM’s pre-trained knowledge and various loss functions to enhance the model’s generalization. Experimental evaluations on diverse datasets demonstrate that AutoMedSAM not only achieves superior segmentation performance but also extends its applicability to clinical environments and non-specialist users. Code is available at https://github.com/HP-ML/AutoPromptMedSAM.git.
@@ 1 Introduction
Deep learning models have traditionally been applied in medicine by designing and training specialized models for specific tasks, achieving significant success [1, 2, 3, 4]. However, these approaches often require training a model from scratch using corresponding data [5, 6, 7], leading to low training efficiency and limited transferability to other tasks [8, 9]. Meanwhile, diagnostic-computer interaction is essential in medical scenarios. Ideally, a physician should be able to provide reference information (e.g., an organ or lesion class ID) to the system, which then recognizes and segments relevant targets. However, most existing specialist models do not support the operation of such interactivity. Recently, prompt-based foundational models in computer vision, such as the Segment Anything Model (SAM) [10], have demonstrated impressive performance and generalization capabilities in various semantic segmentation tasks based on user-provided prompts [11, 12].
@@
In busy clinical environments, it is difficult for physicians to provide accurate prompts through labor-intensive manual prompting, but it is a fundamental requirement for obtaining precise masks. On the other hand, for users without medical knowledge, the pixels in the image corresponding to organs or lesions are not immediately intuitive, and the generated segmentation masks lack additional semantic annotations, making them difficult to interpret. This renders the segmentation results of little practical value for non-expert users, thereby limiting the applicability of MedSAM in certain contexts and its potential user base. Although MedSAM boasts impressive zero-shot and few-shot segmentation capabilities, there are still some challenges. Specifically: (1) Dependency on Manual Prompts: It requires users to provide precise prompts to segment the target region. However, creating prompts for medical images requires expertise, and in clinical settings, providing explicit prompts like points, bounding boxes, or scribbles is impractical. (2) Limitations of Prompt Precision: The manual nature of these prompts means the error may not be within a controllable range [14]. Using the bounding box prompt as an example, MedSAM’s performance relies on the discrepancy between the prompt box and the true boundary. However, other targets within the prompt box cannot be avoided, and different organ or lesion categories often exhibit high similarity and low inter-class variation [15]. (3) Lack of Semantic Information: MedSAM cannot obtain the semantic information of the masks it predicts. It can only predict binary masks for each prompt, without associating them with semantic labels [10, 13, 16].
To address these challenges, we propose AutoMedSAM. A comparison between AutoMedSAM and existing methods is shown in Fig. 1. We replaced the original prompt encoder with a diffusion-based class prompt encoder. This new encoder uses a lesion or organ index as prompting, incorporating semantic information about the target into the AutoMedSAM learning process. The class prompts and image embeddings are input information to generate prompt embeddings for the mask decoder directly. By generating prompt embeddings directly from class prompts, we eliminate the robustness issues caused by manually provided explicit prompts and transform the semi-automatic MedSAM into a fully automated end-to-end process. Meanwhile, we designed an uncertainty-aware joint optimization training strategy. This strategy can optimize the model training process by combining the advantages of multiple loss functions while transferring the pre-training knowledge from MedSAM to AutoMedSAM. This enables AutoMedSAM to adapt to data of various modalities and effectively extract features of different organs or lesions, enhancing the segmentation performance and robustness of the model.
@@
    We introduced an interactive end-to-end model, AutoMedSAM, enabling doctors and researchers to rapidly and accurately analyze medical images for more timely diagnosis and treatment, without relying on labor-intensive manual sketching of organs or lesions to guide the segmentation process.
@@
    We proposed a diffusion-based class prompt encoder that eliminates the need for explicit prompts and facilitates direct learning of latent prompt embeddings from class prompts. Through the diffusion process, the encoder progressively enhances its understanding of the medical targets, thereby improving the effectiveness and generalization capability of the class prompts.
@@
    We designed an uncertainty-aware joint optimization strategy that combines pixel-based, region-based, and distribution-based loss functions. This approach leverages the pre-trained knowledge of MedSAM and integrates diverse loss components to improve the model’s generalization ability and robustness across multi-modal medical data.
@@ 2.1 SAM-based Medical Image Segmentation
SAM represents a significant breakthrough in transforming image segmentation from specialized applications to a general-purpose tool [17]. After training on large-scale datasets, SAM builds a broad knowledge base and relies on manually provided explicit prompts with precise locations (e.g., points and bounding boxes) to trigger segmentation responses [18]. However, due to the substantial domain gap between natural and medical images, SAM exhibits limited generalizability in medical imaging. To bridge this gap, studies such as MedSAM and SAM Med2D integrate extensive medical imaging data with specialized fine-tuning strategies, effectively improving performance in medical scenarios [13, 19, 20]. However, the accuracy of SAM’s segmentation is sensitive to the positional bias of the prompts (as validated in Section 4.5). Consequently, the segmentation process often requires extensive manual intervention or professional detectors, turning it into a multi-stage workflow. In certain clinical settings(e.g., surgeries [15]), providing explicit prompts for every frame is impractical. To address this limitation, methods like MaskSAM and UV-SAM employ task-specific expert models to generate coarse positional prompts, reducing reliance on explicit user input [16, 21]. While these approaches mitigate the need for explicit prompts, they lack fine-grained class-specific prompting capabilities and face scalability challenges [22, 23]. In contrast, our method introduces a class-based prompting approach, leveraging a diffusion process to generate prompt embeddings from image features. This simplified prompting mechanism streamlines the segmentation pipeline, and enhances fine-grained class differentiation.
@@
2.2 Diffusion Models for Medical Domain
Diffusion models show strong potential in medical imaging, achieving notable success in tasks such as image generation, segmentation, and classification [24, 25, 26]. The initial applications of diffusion models in the medical field primarily focused on generating medical data, which has proven useful for medical data augmentation [27]. In addition, several scholars have investigated the potential of medical images generated by diffusion models as a substitute for real data in training deep networks. M. Usman Akbar et al. and D. Stojanovski et al. demonstrated that these synthetic data are effective for downstream tasks [28, 29]. Recent studies have used diffusion models for cross-modality synthesis [30, 31]. For example, DCE-diff addresses data heterogeneity by leveraging multimodal non-contrast images to extract anatomical details from structural MRI sequences and perfusion information from ADC images, enabling the synthesis of early-phase and late-phase DCE-MRI [32]. D. Stojanovski et al. demonstrated that visual realism during model training does not necessarily correlate with model performance. Consequently, these models can generate organ or lesion features optimized for deep learning, thereby enhancing the accuracy of downstream tasks. Furthermore, utilizing more efficient models can significantly reduce computational costs [29]. By progressively refining representations through noise-based generation and denoising, diffusion models inherently capture fine-grained structural details and semantic consistency, which are particularly beneficial for complex medical image tasks.
@@ 3.1 Overview of AutoMedSAM
The optimization of AutoMedSAM is essentially based on the structure of SAM, with the innovative introduction of a diffusion-based class prompt encoder to address the challenges of manual prompts. Specifically, as illustrated in Fig. 2, AutoMedSAM consists of three core modules: an image encoder $E_{I}$, a diffusion-based class prompt encoder $E_{P}$, and a mask decoder $D_{M}$. The input image is denoted as $I\in\mathbb{R}^{h\times w\times 3}$, with spatial resolution $h\times w$. By providing the prompt class $c$, AutoMedSAM can predict the mask corresponding to the class $c$. The image is first processed by the image encoder to generate the image embedding $F_{I}$. Subsequently, the class prompt encoder $E_{P}$, based on a diffusion model, processes the image embedding to generate sparse prompt embedding $P_{s}^{(c)}$ and dense prompt embedding $P_{d}^{(c)}$ from the target prompts. Finally, the mask decoder combines the image embedding, positional encoding $P_{p}$, sparse prompt embedding, and dense prompt embedding to predict the segmentation mask $M^{(c)}$ for class $c$. The entire process can be represented as:
@@ 3.3 Uncertainty-aware Joint Optimization
AutoMedSAM is designed to accommodate various modalities of medical images, which often have significant differences. Moreover, medical images generally have characteristics such as low contrast and high similarity among targets. To enhance segmentation accuracy, the model’s optimization process needs to integrate multiple loss functions to improve overall performance. However, as the number of loss functions increases, adjusting their weights becomes increasingly challenging, and these weights are crucial for model optimization as they directly influence the optimization direction [35]. We introduced an uncertainty-aware joint optimization algorithm to adaptively adjust the loss weights. Without the need for any hyperparameters, this strategy autonomously integrates pixel-level, region-level, distribution-level, and other loss functions to comprehensively enhance the model during training. Specifically, we employed the following loss functions:
@@
(2) Cross-Entropy (CE) Loss: A pixel-based loss that measures the degree of alignment between the predicted class probabilities and the target labels at each pixel.
@@
$L=\sum_{j=1}^{J}\left(\frac{1}{2\lambda_{j}^{2}}L_{j}+\log(1+\lambda_{j}^{2})\right),$ (17)
where $J$ represents the number of loss members involved in the optimization during training, $L_{j}$ represents all the losses we mentioned above, and $\lambda_{j}$ are learnable parameters that adjust the contribution of each loss component based on uncertainty.
@@ 4.1.1 Datasets
To evaluate the generalizability of AutoMedSAM, we conducted tests on several commonly used benchmark datasets, including AbdomenCT1K[37], BraTS[38], Kvasir-SEG[39], and Chest Xray Masks and Labels(Chest-XML)[40]. All data were obtained from the CVPR 2024 Medical Image Segmentation on Laptop Challenge [^1] https://www.codabench.org/competitions/1847. These datasets encompass four distinct imaging modalities: CT, MR, endoscopy, and X-ray, covering diverse segmentation targets ranging from organs to lesions. This diversity effectively demonstrates the broad applicability of our method. Table 1 provides an overview of the datasets used.
@@
AMOS[41] CT(3D) Liver, kidneys, pancreas, spleen
@@ 4.1.5 Baseline Methods
We compared our method with state-of-the-art medical imaging SAM models, grouped into SAM-Core models (SAM [10], SAM2 [42], MedSAM [13], Med2d [20], and U-MedSAM [43]) and SAM-Based models (SAMed [44], H-SAM [45], AutoSAM [46], and SurgicalSAM [15]). The SAM-Core Model retains the foundational framework of SAM, requiring manual prompting for operation. The SAM-Based Model builds upon SAM, introducing enhancements that eliminate the need for manual prompting. Notably, SurgicalSAM, like our approach, uses class prompts and prototype contrastive learning to distinguish surgical instruments. To ensure the reliability of the experimental results, all methods were conducted under identical experimental conditions. Besides, we use expert models (i.e., nnU-Net [47], Swin-Unet [48], and MedFormer [49]) as performance benchmarks.
@@ 4.2 Comparing with the Existing Methods
The interactive segmentation process requires the model to have a multi-objective segmentation capability. Accordingly, we first evaluated the performance of various models on the multi-organ segmentation task using the AbdomenCT1K dataset. The results are shown in Table 2. It can be observed that the performance of different models varies significantly under the DSC and NSD metrics, with AutoMedSAM demonstrating the best overall performance among all methods. Specifically, AutoMedSAM achieves the highest scores on both overall metrics, DSC (94.580%) and NSD (95.148%). This improvement is not only reflected in the average performance, but also stands out at the single-organ level(e.g., Pancreas and Kidney). The performance gain can be attributed to the proposed diffusion-based class prompt encoder and uncertainty-aware joint optimization strategy, which refines the structure and detail of the prompt class through the diffusion process and injects semantic information into the prediction pipeline. In contrast, although SAM-Core models (e.g., MedSAM with an NSD of 92.969%) offer stable performance due to manual prompting, they suffer from inefficiency in practical use. This limitation makes them less suitable for time-sensitive clinical scenarios where fast segmentation of complex medical images is essential. SAM-Based models, while capable of automatic prompting, generally perform worse in terms of segmentation accuracy. All models in this category, except AutoMedSAM, fall below the SAM-Core baseline. Traditional specialist models deliver relatively balanced performance and even excel in specific organs (e.g., nnU-Net with a DSC of 96.317% on the right kidney), yet they lack the generalization and flexibility needed for interactive segmentation tasks, limiting their broader applicability.
Table 2: Comparative Results on the AbdomenCT-1K Dataset. The SAM-Core Model retains the framework of SAM, requiring manual prompting for operation. The SAM-Based Model builds upon SAM, introducing enhancements that eliminate the need for manual prompting. L. and R. stand for left and right. $\uparrow$ means higher is better. The best results are shown in Bold.
Method Category Method DSC(%$\uparrow$ NSD(%$\uparrow$ Organ
DSC(%$\uparrow$ NSD(%$\uparrow$
Specialist Model nnU-Net [47] 93.196 92.621 96.187 96.130 82.095 95.249 96.317 90.962 97.573 83.351 94.103 97.118
Swin-Unet [48] 87.996 86.215 97.057 94.790 68.878 88.874 90.382 96.132 97.126 65.835 80.419 91.563
MedFormer [49] 92.479 88.452 96.883 94.816 80.942 94.861 94.894 95.491 90.733 85.066 87.668 83.300
SAM-Core Model SAM[10] 89.790 83.940 92.117 93.766 72.694 95.254 95.121 75.376 90.253 69.687 92.385 91.998
SAM2[42] 90.191 85.137 93.815 95.391 70.793 95.623 95.334 81.804 94.149 62.519 93.860 93.354
MedSAM[13] 93.505 92.969 96.836 97.120 81.648 96.121 95.800 91.908 98.807 83.845 95.374 94.913
Med2d[20] 83.840 79.347 93.562 93.371 58.168 87.081 87.020 80.020 89.751 67.776 79.624 79.566
SAM-Based model SAMed[44] 81.329 78.504 97.132 96.838 76.512 67.904 68.259 92.435 95.390 77.635 63.342 63.716
AutoMedSAM(Ours) 94.580 95.148 97.467 96.958 86.061 96.291 96.121 95.030 98.911 88.903 96.585 96.309
@@
As shown in Table 2, the segmentation performance of the tested models varies minimally for the liver, spleen, and pancreas but significantly deteriorates for kidney segmentation. To investigate this, we visualized the segmentation results (shown in Fig. 4). Fig. 4 (a) and (b) illustrate that while SAM-Based models generally perform well in boundary recognition, they confuse the left and right kidneys. This is due to the strong symmetry and high morphological similarity of the left and right kidneys, and the SAM-Core model addresses this challenge by providing location information of the target organs through manual prompting. However, the automatic prompting mechanism of the SAM-Based model lacks this information. This is why their performance drops sharply. However, during manual prompting, bounding boxes may unavoidably include other organs. Since SAM-Core models lack semantic information during prediction, they cannot accurately identify specific organs, leading to segmentation errors. As shown in Fig. 4 (c) and (d), both SAM and MedSAM mistakenly identified other tissues to varying degrees. Additionally, even with minimal redundancy within bounding boxes, SAM models tend to segment features that appear prominent, which degrades mask quality (shown in Fig. 4 (e) and (f)). In contrast, our method incorporates class-based prompts to introduce semantic information into predictions, effectively mitigating this issue.
@@
Figure 5: The qualitative analysis results of AutoMedSAM and other comparison models on BraTS, Kvasir-SEG, and Chest-XML.
To evaluate segmentation accuracy independently, we conducted additional experiments on the BraTS, Kvasir-SEG, and Chest-XML datasets. In BraTS and Kvasir-SEG, the targets (tumor and polyp) are singular but structurally complex, with diverse shapes and ambiguous boundaries. In Chest-XML, overlapping structures like ribs around the lungs and the large target area introduce significant challenges. These datasets provide diverse difficulties suitable for evaluating segmentation performance. The results and visualizations are shown in Table 3 and Fig. 5, respectively.
Table 3: Comparative Results on the BraTS, Kvasir-SEG and Chest-XML Datasets
@@
BraTS(Tumor) Kvasir-SEG(Polyp) Chest-XML(Lung)
@@
U-MedSAM[43] 89.616 88.793 95.007 97.091 96.813 97.310
SAM-Based model SAMed[44] 89.902 89.097 86.393 88.255 94.066 94.600
H-SAM[45] 90.615 90.288 88.010 90.084 92.264 92.978
AutoSAM[46] 90.240 89.186 87.045 89.288 89.616 90.516
@@
AutoMedSAM(Ours) 91.057 92.661 96.828 98.729 96.941 97.367
As shown in Table 3, AutoMedSAM achieved superior performance across all tasks, with DSC and NSD scores of 96.828% and 98.729% for polyp segmentation, demonstrating its adaptability to complex medical environments. Conversely, when SAM-Core models lose their advantage of manual prompts, their performance declines significantly. For example, in BraTS, SAM achieved only 69.667% DSC and 42.112% NSD for tumor segmentation, indicating reasonable overlap with ground truth but poor boundary recognition. Similar issues were observed in lung segmentation, as shown in the last two subfigures of Fig. 5, where SAM models produced masks containing excessive non-target tissues. In conclusion, the experimental results confirm that the proposed AutoMedSAM effectively adapts to various medical modalities, delivering accurate segmentation masks and achieving the best metrics across all datasets.
4.3 Cross-Dataset Generalization
To verify the cross-dataset and cross-modality generalization of AutoMedSAM, we trained the model on AbdomenCT-1K and BraTS-FLAIR, then tested it on AMOS and BraTS-T1CE. The results are shown in Table 4, where only the organ or lesion classes shared by both datasets are considered. Our method holds significant advantages compared to SurgicalSAM, which also employs class indexes as prompts. Notably, when generalizing from BraTS-FlAIR to BraTS-T1CE, we achieve a large improvement of 15.67% in the DSC. This demonstrates that, compared to SurgicalSAM’s prototype-based contrastive learning, AutoMedSAM achieves stronger generalization ability by capturing more fine-grained features of organs or lesions through the diffusion process.
Table 4: Cross-Dataset Generalization.
Training Testing Method DSC(%$\uparrow$ NSD(%$\uparrow$
AbdomenCT-1K AMOS SurgicalSAM 56.930 64.451
AutoMedSAM 71.141 77.282
BraTS-FLAIR BraTS-T1CE SurgicalSAM 36.843 39.825
AutoMedSAM 52.513 48.101